11 A learner in your class says that the of the underlined digit in 12 678 is 2. How will you help the learner to overcome this misconception? (2)


12 What does partitioning mean? Explain and illustrate your answer. (3)


13 In the context of choosing a "whole", explain when a "quarter" is not always equal to a "quarter". Give an example. (2)


14 Use the correct language to explain the difference between the expressions:

"It is the fifth of December"

"You are the fifth in the row"

"I want a fifth of the pizza"

What types of numbers are we dealing with in these expressions? (4)


1
Expert's answer
2021-05-26T16:37:34-0400

11.

The teacher should develop pretest misconceptions based on their knowledge of the course.

They should then focus on explaining why right is right before addressing the misconceptions

The teacher then should activate the misconception but minimize focus on the misconception and finally, they should revisit the facts why right is right.


12.

Partitioning is a method of solving maths problems involving large numbers by first splitting them into smaller units so that they can be easily computed.

For example,

79+34=113


The minors are taught to split it into

70+9+30+4=113

Or 70+30+9+4=113

13.

Quarter of a whole is not equal to quarter when the whole is not equal to 1. For example, If the whole is 2, Quarter will be 0.5.


14.

"It is the fifth of December"

The statement answers the question when. It tells of the current date as being December 5


"You are the fifth in the row"

The statement tells the person's position as being position 5 in a row


"I want a fifth of the pizza"

The statement refers to a fraction of pizza. The person wants to be offered 1/5 fraction of the Pizza.